OpenStudy (anonymous):

help fast will fan/metal solve the inequality x over -10>6 x=? 5z is greater than or equal to -75 z=?

OpenStudy (ahsome):

Is it: \[\frac{x}{-10}<6\]

OpenStudy (anonymous):

\[\frac{ x }{ 10}>6\]

OpenStudy (ahsome):

Not \(-10\)?

OpenStudy (anonymous):

its -10 srry

OpenStudy (ahsome):

Ok, so this is it: \[\frac{x}{-10}<6\] Now, we need to multiply both sides with \(-10\) to leave \(x\) by itself \[\frac{x}{-10}\times-10<6\times-10\] When we multiply or divide any number that is negative, the greater or less sign flips. So now, it is: \[\frac{x}{-10}\times-10>6\times-10\] Does that make sense?

OpenStudy (anonymous):

little more explanation plz

OpenStudy (ahsome):

Ok, we have the equation: \[\frac{x}{-10}>6\] Now, we want to keep the \(x\) by itself. How do we do this?

OpenStudy (anonymous):

x=? A.x>60 B. x>-60 C. x<-60 D. x=-4